# AIRMap — AI-Generated Radio Maps

### v1

@@ V Radio Maps for Real-Time Digital Twins
As discussed in Section I, the ultimate objective of these radio maps is to enable real‑time estimation of channel parameters across a geographic area. The next step is to integrate our U‑Net model into higher‑layer workflows for system‑level evaluation. In this section, we demonstrate and evaluate how the U‑Net serves as a channel‑parameter generator for both Sionna SYS system simulations [19], and Colosseum—the world’s largest wireless network emulator [54].
@@ V-A Evaluation on Sionna SYS
We evaluate how path gain prediction errors impact system-level performance, focusing on spectral efficiency and Block Error Rate (BLER), as well as the feasibility of real-time simulation. For this analysis, we use the Sionna SYS simulator [19], which enables realistic wireless network simulations through a flexible and Python-based interface.
Fig. 15 compares errors for three channel‑modeling approaches—Sionna RT 0.19.2, the uncalibrated U‑Net, and the calibrated U‑Net—against measurement data. Each model employs a single‑tap channel with fixed propagation delay and randomized phase to focus exclusively on path‑gain effects in system‑level evaluations.

### v3

@@ V Radio Maps for Real-Time Digital Twins
As introduced in Section I, the goal of these radio maps is to enable real-time estimation of channel parameters over a geographic area for wireless digital twins—real-time, scalable simulators that replicate the behavior of actual wireless networks. The next step is to integrate our U-Net model into higher-level workflows for system-level evaluation. In this section, we demonstrate and evaluate how the U-Net functions as a channel-parameter generator for both Sionna SYS system simulations [20] and Colosseum—the world’s largest wireless network emulator [36]. To ensure a consistent and fair comparison, all system-level experiments were conducted using the same trajectory and measurement described in Section II, enabling direct evaluation of AIRMap’s performance against empirical ground-truth data.
@@ V-A Evaluation on Sionna SYS
We evaluate how path gain prediction errors impact system-level performance, focusing on spectral efficiency and BLER, as well as the feasibility of real-time simulation. For this analysis, we use the Sionna SYS simulator [20], which enables realistic wireless network simulations through a flexible and Python-based interface.
Fig. 17 compares errors for three channel‑modeling approaches—Sionna RT, the uncalibrated U‑Net, and the calibrated U‑Net—against measurement data. Each model employs a single‑tap channel with fixed propagation delay and randomized phase to focus exclusively on path‑gain effects in system‑level evaluations.